OpenStudy (anonymous):

Two marbles are drawn without replacement from a box with 3 white, 2 green, and 1 blue marble. Find the probability of drawing one white marble and one blue marble.

OpenStudy (raden):

P(white) = 3/6 = 1/2 P(blue) = 1/5 then multiply both values above: 1/2 * 1/5 = ... ? (this is as the answer)